# Stable Trajectory Clustering: An Efficient Split and Merge Algorithm

### v1

@@ III Preliminaries and Definitions
Definition III.2 (Trajectory line segment).
@@
Definition III.4 (Average Euclidean distance of line segments).
@@
[…]
$\begin{cases}x_{k}\left(t_{\sigma}\right)=x_{k}\left(t_{i+1}\right)\times t_{\sigma}+x_{k}\left(t_{i}\right)\times\left(1-t_{\sigma}\right),\\ y_{k}\left(t_{\sigma}\right)=y_{k}\left(t_{i+1}\right)\times t_{\sigma}+y_{k}\left(t_{i}\right)\times\left(1-t_{\sigma}\right).\end{cases}$ (2)
Lemma III.1.
@@
$dist\left(P_{k}^{t_{\sigma}},P_{j}^{t_{\sigma}}\right)$ $=\sqrt{\left(x_{k}\left(t_{\sigma}\right)-x_{j}\left(t_{\sigma}\right)\right)^{2}+\left(y_{k}\left(t_{\sigma}\right)-y_{j}\left(t_{\sigma}\right)\right)^{2}}.$ (3)
@@
$=\frac{1}{\tau}\evaluated{\left(\frac{(2at_{\sigma}+b)\sqrt{at_{\sigma}^{2}+bt_{\sigma}+c}}{4a}\right)}_{t_{i}}^{t_{i+1}}$
$-\evaluated{\frac{1}{\tau}\left(\frac{(b^{2}-4ac)\ln(2\sqrt{a}\sqrt{at_{\sigma}^{2}+bt_{\sigma}+c}+2at_{\sigma}+b)}{8a^{\frac{3}{2}}}\right)}_{t_{i}}^{t_{i+1}}.$
@@
Definition III.5 ($\varepsilon$-neighborhood).
@@
Definition III.6 (Core line segment).
@@
Definition III.7 (Directly density-reachable).
@@
Definition III.8 (Density-reachable).
@@
Definition III.9 (Density-connected).
@@
Definition III.10 (Borderline segment).
@@
Definition III.11 (Outlier line segment).
@@
Definition III.12 (Cluster of line segments).
@@
IV Clustering Scheme
As mentioned in previous clustering studies, the split and merge operations are two significant events for moving object clustering  [29, 27]. In this paper, we also support this view of clustering line segments in each time interval. The split event for line segments involves identifying which line segments are separated from each other, while the merge event involves combining line segments that are close to each other and have similar properties.
@@ 19 $C_{s}\leftarrow$ subTraj($C_{hist},m,\Pi,S,W$);
The details of this algorithm are as follows: Firstly, the preprocessing should be run on data trajectories to divide each trajectory into $m$ parts (lines 1 and 2). Then, in the first time interval, for obtaining the first clusters, DBSCAN line segment clustering should be executed (line 3). From the second time interval to the last time interval (lines 5-16), if any cluster of line segments exists, the Split function (Algorithm IV-A) is called (line 8). Two outputs are obtained: either the segments remain in their cluster or they are split from each other. The Merge function (Algorithm 3) should then be executed based on these results (lines 9-15). Eventually, the results of line segment clustering are utilized as inputs for the whole-trajectory clustering (Algorithm 4) and the sub-trajectory clustering (Algorithm 5) algorithms (lines 17 and 18).
IV-A Split of Line Segments
According to definition III.12, let $C_{s}$ be a cluster containing $n_{s}$ line segments in the time interval $\left[t_{i-1},t_{i}\right]$, where $n_{s}\leq n$. In the following, we will provide a definition and explanations regarding the split of the cluster.
Definition IV.1 (Split event).
@@ IV-A Split of Line Segments
Suppose that $UC_{s}$ is a set of line segments in the time interval $\left[t_{i},t_{i+1}\right]$ that have previously been in cluster $C_{s}$. To detect whether the cluster $C_{s}$ will be split in the next time interval or not, a random line segment in $UC_{s}$ should be selected initially and then its neighboring line segments are determined based on their average Euclidean distance as sub-set $c_{s_{1}}$. The same process is repeated for the remaining line segments in $UC_{s}$ until all line segments are examined, and no line segments are remaining. So, there is a set of sub-sets $NC_{s}=\left\{c_{s_{1}},c_{s_{2}},...,c_{s_{num}}\right\}$ and we need to determine whether the members of each sub-set can be considered neighbors of each other or not. If there is no chain of line segments that connects two sub-sets, they are considered split from each other; otherwise, they are connected. So, the $NC_{s}$ should be updated such that its members are the new sub-sets where, according to definition III.12, each sub-set satisfies the Maximality criterion. However, for each subset to be a cluster, we also need to check for Connectivity. Consequently, we have sub-sets that are either considered as a sufficiently dense cluster or are recognized only as a low-density cluster of $NC_{s}$ due to violating the Connectivity criterion. It is noteworthy that it is possible to preprocess the line segments before using formula (III.4), meaning we can measure only the Euclidean distance between the start and end points of the line segments without using the average Euclidean distance of line segments. Therefore, the following cases occur between the start and end points of each pair of line segments in $UC_{s}$ at the time interval $\left[t_{i},t_{i+1}\right]$:
@@
Algorithm IV-A provides the pseudocode for the split event used in trajectory clustering. The algorithm takes a set of densely clustered line segments in a time interval and determines whether these line segments remain connected or are split from each other in the next time interval. If they are split, the resulting sub-clusters are then evaluated to check if they are dense enough. This involves checking the distances between line segments to identify neighbors, ensuring maximal connectivity between clusters, and assessing the density of the clusters based on the presence of core line segments.
input : Line segments set $UC_{s}=\left\{l_{1},l_{2},...,l_{n_{s}}\right\}$ in time interval $\left[t_{i-1},t_{i}\right]$, $\varepsilon$, $MinLns$.
@@ 30      $c_{s_{i}}$ is a dense enough cluster;
31      $c_{s_{i}}$ is a outlier or low-density cluster;
@@
return $NC_{s}$  The detailed steps of this algorithm are as follows: the algorithm finds the neighboring line segment for each line segment in $UC_{s}$ and creates clusters without determining whether they are sufficiently dense or not (lines 2-7). The algorithm then checks the Maximality (lines 9-19) and Connectivity (lines 21-25) criteria, as mentioned in definition III.12.
IV-B Merge of Line Segments
@@
Definition IV.2 (Merge event).
@@
To identify the two line segments that lead to the merge of two clusters, a simple search is conducted between the elements of the two clusters. The search stops as soon as two line segments are found that have an average Euclidean distance of less than or equal to $\varepsilon$. The merge event may decrease the number of clusters, however, and determining whether new clusters have the required density depends on the user input parameters set for the DBSCAN clustering algorithm. Algorithm 3 merges two clusters of line segments based on a proximity threshold and evaluates whether the resulting cluster is dense. The density is determined by the presence of core line segments and the total number of line segments in the merged cluster.
input : Two non-empty low-density or dense enough clusters $c_{m_{1}}$ and $c_{m_{2}}$, $\varepsilon$, $MinLns$.
@@
1 foreach line segment $l_{i}$ in $c_{m_{1}}$ do foreach line segment $l_{j}$ in $c_{m_{2}}$ do if $dist(l_{i},l_{j})\leq\varepsilon$ then $c_{m_{1}m_{2}}\leftarrow c_{m_{1}}\cup c_{m_{2}}$;
@@
7 if $|c_{m_{1}m_{2}}|=|c_{m_{1}}|+|c_{m_{2}}|$ then if $c_{m_{1}m_{2}}$ contains at least one core line segment  then $c_{m_{1}m_{2}}$ is a dense enough cluster; else $c_{m_{1}m_{2}}$ is a low-density cluster; end if
8return $c_{m_{1}m_{2}};$ else return $\emptyset$;
9 end if
@@
In this algorithm, two sufficiently dense or two non-empty low-density clusters are given as inputs. Then it is checked which two line segments can connect two clusters (lines 1-6). If two clusters are connected, then the Connectivity should be checked (lines 7-14). It is noteworthy that in some time intervals, both split and merge events occur. The Split event never executes on the outlier line segments. Instead, only the Merge event is applied to them, which alters the role of the line segments in the clusters they join.
IV-C Whole-trajectory Clustering and Sub-trajectory Clustering
In whole-trajectory clustering, the algorithm considers the entire trajectory of an entity. In other words, features like average Euclidean distance and constant velocity of moving objects in each time interval hold significance for clustering. With this in mind, we can use the clustered line segment information obtained by the split and merge events which we discussed in previous sections. Thus, based on definition III, two whole trajectories $\pi_{i}$ and $\pi_{j}$ can be assigned to the same cluster if their line segments consistently belong to the same cluster; either they are neighbors of each other directly or indirectly across all time intervals (definitions III.8 and III.7). Algorithm 4 provides whole-trajectory clustering by comparing trajectories at multiple time intervals to identify similarities. It counts the number of time intervals during which the line segments of each pair of trajectories are in the same cluster. If this count matches the total number of time intervals, the trajectories in the pair are considered similar because they consistently follow similar paths over time. The algorithm constructs a joint or disjoint graph that represents these similar relationships between trajectories and returns each connected component in this graph as a cluster of similar trajectories.
@@ 1 $\mathcal{A}_{\mathtt{traj}}\leftarrow\mathbb{O}_{m\times m};$
2 foreach trajectory $\pi_{i}\in\Pi$  do $counter\leftarrow 0;$
3 foreach trajectory $\pi_{j}\in\Pi$ where $j>i$ do for t from 1 to $m$ do for k from 1 to $|C_{hist}\{t\}|$ do if $l_{\pi_{i}}$ and $l_{\pi_{j}}\in C_{hist}\{t\}\{k\}$  then $counter\leftarrow counter+1;$ end if
4end for
5end for
[…]
7end foreach
@@
Algorithm 4 Whole-trajectory clusteirng
In this algorithm, a matrix is used to determine which whole trajectories are in the same cluster in all time intervals (lines 4-9). Subsequently, the graph structure is employed to determine the whole-trajectory clusters. Each graph represents a cluster of whole trajectories. Besides that, a node corresponds to a trajectory, and the edge between two nodes signifies their neighborhood (line 10). For sub-trajectory clustering, we use the sliding window model. This model is appropriate when we aim to detect specific ranges within a given dataset and assess the patterns, trends, or deviations within those ranges  [45, 46, 47]. As mentioned in the definition III, the goal of sub-trajectory clustering is to recognize the time interval ranges like $R_{p}$ and cluster the sub-trajectories in this range as $C_{p}^{R_{p}}$. Utilizing the sliding window model as a tool can facilitate the process of finding these time interval ranges more easily. Therefore, we should first set the size of the sliding window and identify how many steps we want to move it. We consider two parameters $W$ and $S$ as the sliding window size and the movement step forward, respectively, which depend on the time interval range. To give an illustration, if each time interval range is considered 5 seconds, for $W=3$ and $S=1$, the sliding window encompasses the data of 3 consecutive time intervals, which are 15 seconds, and moves forward 5 seconds for each step (Fig. 2(e)). The window moves across all time intervals and according to the window size, contains the line segment clusters obtained in the previous section by the split and the merge events. In the current window, the consecutive time intervals that incorporate similar line segment clusters are determined as sub-trajectory clusters. Based on the parameter $S$, moving the sliding window either combines similar sub-trajectory clusters with the previous ones or creates new sub-trajectory clusters. If $S<W$, then the consecutive windows contain common time intervals and line segment clusters. Algorithm 5 describes sub-trajectory clustering. It processes pre-clustered line segments to identify and cluster similar sub-trajectories. It employs a sliding window approach, iterating over the line segments according to parameters $S$ and $W$. For each window, sub-trajectories are extracted and clustered, with the results saved incrementally. After processing all windows, consecutive clustering results are compared to identify similar and different ranges. Continuous ranges of similar results are merged, while different ranges are noted separately. The algorithm then compiles these identified ranges into a final list of clustered sub-trajectories, resulting in a set of clusters that represent similar segments within the overall trajectory data.
@@
1 for i from 1 to $|C_{hist}|$+W-1 by step S do $W_{s}\leftarrow i;$ $W_{e}\leftarrow i+W-1;$

### v2

@@ 3. Preliminaries and Definitions
Definition 3.0 (Trajectory line segment).
@@
Although trajectory dissimilarity can be measured using various metrics (e.g., Fréchet distance, DTW (58)), we adopt the average Euclidean distance because it is the most appropriate choice for piecewise-linear trajectories and fits our clustering framework, rather than being selected merely for convenience or because our algorithm only works with this metric. The distance is defined as follows:
Definition 3.0 (Average Euclidean distance of line segments).
@@
(1) $dist(l_{k},l_{j})=\frac{1}{\tau}\int_{\tau}\|P^{t}_{k}-P^{t}_{j}\|\,dt,$
where $P^{t}_{k}=(x_{k}(t),y_{k}(t))$ and $P^{t}_{j}=(x_{j}(t),y_{j}(t))$ represent the positions of the objects at time $t$ on the line segments $l_{k}$ and $l_{j}$, respectively. The notation $\|\cdot\|$ denotes the Euclidean distance between the points at time $t$, and $\tau=t_{i+1}-t_{i}$ represents the time interval. For any arbitrary $t_{\sigma}\in[0,1]$, there is a geographical position $P_{k}^{t_{\sigma}}$ on line segment $l_{k}$, which is defined by the convex set as follows. This formula also holds for $P_{j}^{t_{\sigma}}$.
(2) $\begin{cases}x_{k}\left(t_{\sigma}\right)=x_{k}\left(t_{i+1}\right)\times t_{\sigma}+x_{k}\left(t_{i}\right)\times\left(1-t_{\sigma}\right),\\ y_{k}\left(t_{\sigma}\right)=y_{k}\left(t_{i+1}\right)\times t_{\sigma}+y_{k}\left(t_{i}\right)\times\left(1-t_{\sigma}\right).\end{cases}$
Lemma 3.5.
@@
(3) $dist\left(P_{k}^{t_{\sigma}},P_{j}^{t_{\sigma}}\right)$ $=\sqrt{\left(x_{k}\left(t_{\sigma}\right)-x_{j}\left(t_{\sigma}\right)\right)^{2}+\left(y_{k}\left(t_{\sigma}\right)-y_{j}\left(t_{\sigma}\right)\right)^{2}}.$
@@
(4) $dist(l_{k},l_{j})=\frac{1}{\tau}\int_{t_{i}}^{t_{i+1}}\sqrt{\left(A^{2}+C^{2}\right)t_{\sigma}^{2}+\left(2AB+2CD\right)t_{\sigma}+B^{2}+D^{2}}\,dt_{\sigma}.$
If we set $a=A^{2}+C^{2},\quad b=2AB+2CD,\quad c=B^{2}+D^{2}$, then
$dist(l_{k},l_{j})$ $=\frac{1}{\tau}\int_{t_{i}}^{t_{i+1}}\sqrt{at_{\sigma}^{2}+bt_{\sigma}+c}\,dt_{\sigma}$
@@
Definition 3.0 ($\varepsilon$-neighborhood).
@@
Definition 3.0 (Core line segment).
@@
Definition 3.0 (Directly density-reachable).
@@
Definition 3.0 (Density-reachable).
@@
Definition 3.0 (Density-connected).
@@
Definition 3.0 (Borderline segment).
@@
Definition 3.0 (Outlier line segment).
@@
Definition 3.0 (Cluster of line segments).
@@
4. Clustering Scheme
As mentioned in previous clustering studies, the split and merge operations are two significant events for moving object clustering  (9; 11). In this paper, we also support this view of clustering line segments in each time interval. The split event for line segments involves identifying which line segments are separated from each other, while the merge event involves combining line segments that are close to each other and have similar properties.
@@ 6 $C_{new}\leftarrow\emptyset;$
12       $C_{mrg}\leftarrowMerge$$(c_{i},c_{j},\varepsilon,MinLns));$
13       if $C_{mrg}\neq\emptyset$ then
14           remove $c_{i}$ and $c_{j}$;
@@ 19 $C_{s}\leftarrow$ subTraj($C_{hist},m,\Pi,S,W$);
The details of this algorithm are as follows: Firstly, the preprocessing should be run on data trajectories to divide each trajectory into $m$ parts (lines 1 and 2). Then, in the first time interval, for obtaining the first clusters, DBSCAN line segment clustering should be executed (line 3). From the second time interval to the last time interval (lines 5-16), if any cluster of line segments exists, the Split function (Algorithm 2) is called (line 8). Two outputs are obtained: either the segments remain in their cluster, or they are split from each other. The Merge function (Algorithm 3) should then be executed based on these results (lines 9-15). Eventually, the results of line segment clustering are utilized as inputs for the whole-trajectory clustering (Algorithm 4) and the sub-trajectory clustering (Algorithm 5) algorithms (lines 17 and 18). Here, in the first time interval, the initialization cost is $\mathcal{O}(n^{2})$ (or $\mathcal{O}(n\log n)$ with spatial indexing). For the remaining $m$ time intervals, let $\mathcal{C}_{split}$ and $\mathcal{C}_{merge}$ denote the computational costs of the split and merge operations, respectively, and let $k$ be the number of clusters in each interval. Thus, the total worst-case complexity for Split-Merge part (lines 1-16) is $\mathcal{O}(n^{2}+m(k\cdot\mathcal{C}_{split}+k^{2}\cdot\mathcal{C}_{merge}))$.
4.1. Split of Line Segments
According to definition 3.13, let $C_{s}$ be a cluster containing $n_{s}$ line segments in the time interval $\left[t_{i-1},t_{i}\right]$, where $n_{s}\leq n$. In the following, we will provide a definition and explanations regarding the split of the cluster.
Definition 4.0 (Split event).
@@ 4.1. Split of Line Segments
Suppose that $UC_{s}$ is a set of line segments in the time interval $\left[t_{i},t_{i+1}\right]$ that have previously been in cluster $C_{s}$. To detect whether the cluster $C_{s}$ will be split in the next time interval or not, a random line segment in $UC_{s}$ should be selected initially, and then its neighboring line segments are determined based on their average Euclidean distance as sub-set $c_{s_{1}}$. The same process is repeated for the remaining line segments in $UC_{s}$ until all line segments are examined, and no line segments are remaining. So, there is a set of sub-sets $NC_{s}=\left\{c_{s_{1}},c_{s_{2}},...,c_{s_{num}}\right\}$ and we need to determine whether the members of each sub-set can be considered neighbors of each other or not. If there is no chain of line segments that connects two sub-sets, they are considered split from each other; otherwise, they are connected. So, the $NC_{s}$ should be updated such that its members are the new sub-sets where, according to definition 3.13, each sub-set satisfies the Maximality criterion. However, for each subset to be a cluster, we also need to check for Connectivity. Consequently, we have sub-sets that are either considered as a sufficiently dense cluster or are recognized only as a low-density cluster of $NC_{s}$ due to violating the Connectivity criterion. It is noteworthy that it is possible to preprocess the line segments before using formula (3.4), meaning we can measure only the Euclidean distance between the start and end points of the line segments without using the average Euclidean distance of line segments. Therefore, the following cases occur between the start and end points of each pair of line segments in $UC_{s}$ at the time interval $\left[t_{i},t_{i+1}\right]$:
@@
Algorithm 2 provides the pseudocode for the split event used in trajectory clustering. The algorithm takes a set of densely clustered line segments in a time interval and determines whether these line segments remain connected or are split from each other in the next time interval. If they are split, the resulting sub-clusters are then evaluated to check if they are dense enough. This involves checking the distances between line segments to identify neighbors, ensuring maximal connectivity between clusters, and assessing the density of the clusters based on the presence of core line segments.
input : Line segments set $UC_{s}=\left\{l_{1},l_{2},...,l_{n_{s}}\right\}$ in time interval $\left[t_{i-1},t_{i}\right]$, $\varepsilon$, MinLnsMinLns.
@@ 30 $c_{s_{i}}$ is a dense enough cluster;
31 $c_{s_{i}}$ is an outlier or low-density cluster;
32 return $NC_{s}$;
@@ 31 $c_{s_{i}}$ is an outlier or low-density cluster;
The detailed steps of this algorithm are as follows: the algorithm finds the neighboring line segment for each line segment in $UC_{s}$ and creates clusters without determining whether they are sufficiently dense or not (lines 2-7). The algorithm then checks the Maximality (lines 9-19) and Connectivity (lines 21-25) criteria, as mentioned in definition 3.13. The time complexity for checking the distances between all line segment pairs in $UC_{s}$ is $\mathcal{O}(|UC_{s}|^{2})$.
4.2. Merge of Line Segments
@@
Definition 4.0 (Merge event).
@@
To identify the two line segments that lead to the merge of two clusters, a simple search is conducted between the elements of the two clusters. The search stops as soon as two line segments are found that have an average Euclidean distance of less than or equal to $\varepsilon$. The merge event may decrease the number of clusters, however, and determining whether new clusters have the required density depends on the user input parameters set for the DBSCAN clustering algorithm.
Algorithm 3 merges two clusters of line segments based on a proximity threshold and evaluates whether the resulting cluster is dense. The density is determined by the presence of core line segments and the total number of line segments in the merged cluster.
input : Two non-empty low-density or dense enough clusters $c_{m_{1}}$ and $c_{m_{2}}$, $\varepsilon$, MinLnsMinLns.
@@
1 foreach line segment $l_{i}$ in $c_{m_{1}}$ do
2 foreach line segment $l_{j}$ in $c_{m_{2}}$ do
3    if $dist(l_{i},l_{j})\leq\varepsilon$ then
4       $c_{m_{1}m_{2}}\leftarrow c_{m_{1}}\cup c_{m_{2}}$;
@@
7 if $|c_{m_{1}m_{2}}|=|c_{m_{1}}|+|c_{m_{2}}|$ then
8 if $c_{m_{1}m_{2}}$ contains at least one core line segment then
9    $c_{m_{1}m_{2}}$ is a dense enough cluster;
10 else
11    $c_{m_{1}m_{2}}$ is a low-density cluster;
12 return $c_{m_{1}m_{2}};$
13else
14 return $\emptyset$;
@@ 11    $c_{m_{1}m_{2}}$ is a low-density cluster;
In this algorithm, two sufficiently dense or two non-empty low-density clusters are given as inputs. Then it is checked which two line segments can connect two clusters (lines 1-6). If two clusters are connected, then the Connectivity should be checked (lines 7-14). It is noteworthy that in some time intervals, both split and merge events occur. The Split event never executes on the outlier line segments. Instead, only the Merge event is applied to them, which alters the role of the line segments in the clusters they join. In the worst case, the algorithm performs $\mathcal{O}(|cm_{1}|\cdot|cm_{2}|)$ comparisons.
4.3. Whole-trajectory Clustering and Sub-trajectory Clustering
In whole-trajectory clustering, the algorithm considers the entire trajectory of an entity. In other words, features like average Euclidean distance and constant velocity of moving objects in each time interval hold significance for clustering. With this in mind, we can use the clustered line segment information obtained by the split and merge events which we discussed in previous sections. Thus, based on definition 3, two whole trajectories $\pi_{i}$ and $\pi_{j}$ can be assigned to the same cluster if their line segments consistently belong to the same cluster; either they are neighbors of each other directly or indirectly across all time intervals (definitions 3.9 and 3.8).
Algorithm 4 provides whole-trajectory clustering by comparing trajectories at multiple time intervals to identify similarities. It counts the number of time intervals during which the line segments of each pair of trajectories are in the same cluster. If this count matches the total number of time intervals, the trajectories in the pair are considered similar because they consistently follow similar paths over time. The algorithm constructs a joint or disjoint graph that represents these similar relationships between trajectories and returns each connected component in this graph as a cluster of similar trajectories.
@@ 1 $\mathcal{A}_{\mathtt{traj}}\leftarrow\mathbb{O}_{m\times m};$
2 foreach trajectory $\pi_{i}\in\Pi$ do
3 $counter\leftarrow 0;$
4 foreach trajectory $\pi_{j}\in\Pi$ where $j>i$ do
5    for t from 1 to $m$ do
6       for k from 1 to $|C_{hist}\{t\}|$ do
7          if $l_{\pi_{i}}$ and $l_{\pi_{j}}\in C_{hist}\{t\}\{k\}$ then
8             $counter\leftarrow counter+1;$
9    $\mathcal{A}_{\mathtt{traj}}(i,j)\leftarrow(counter=m)\;?\;1:0;$
@@
Algorithm 4 Whole-trajectory clustering
In this algorithm, a matrix is used to determine which whole trajectories are in the same cluster in all time intervals (lines 4-9). Subsequently, the graph structure is employed to determine the whole-trajectory clusters. Each graph represents a cluster of whole trajectories. Besides that, a node corresponds to a trajectory, and the edge between two nodes signifies their neighborhood (line 10). The time complexity of the algorithm is $\mathcal{O}(n^{2}\cdot m)$, because it compares the histories of line segment clusters for every trajectory pair over $m$ time intervals. Additionally, standard DFS or BFS computes the connected components of the resulting graph in $\mathcal{O}(n+n^{2})$ time.
For sub-trajectory clustering, we use the sliding window model. This model is appropriate when we aim to detect specific ranges within a given dataset and assess the patterns, trends, or deviations within those ranges  (41; 42; 43). As mentioned in the definition 3.3, the goal of sub-trajectory clustering is to recognize the time interval ranges like $R_{p}$ and cluster the sub-trajectories in this range as $C_{p}^{R_{p}}$. Utilizing the sliding window model as a tool can facilitate the process of finding these time interval ranges more easily. Therefore, we should first set the size of the sliding window and identify how many steps we want to move it. We consider two parameters $W$ and $S$ as the sliding window size and the movement step forward, respectively, which depend on the time interval range. To give an illustration, if each time interval range is considered 5 seconds, for $W=3$ and $S=1$, the sliding window encompasses the data of 3 consecutive time intervals, which are 15 seconds, and moves forward 5 seconds for each step (Fig. 2(e)). The window moves across all time intervals and according to the window size, contains the line segment clusters obtained in the previous section by the split and the merge events. In the current window, the consecutive time intervals that incorporate similar line segment clusters are determined as sub-trajectory clusters. Based on the parameter $S$, moving the sliding window either combines similar sub-trajectory clusters with the previous ones or creates new sub-trajectory clusters. If $S<W$, then the consecutive windows contain common time intervals and line segment clusters.
Algorithm 5 describes sub-trajectory clustering. It processes pre-clustered line segments to identify and cluster similar sub-trajectories. It employs a sliding window approach, iterating over the line segments according to parameters $S$ and $W$. For each window, sub-trajectories are extracted and clustered, with the results saved incrementally. After processing all windows, consecutive clustering results are compared to identify similar and different ranges. Continuous ranges of similar results are merged, while different ranges are noted separately. The algorithm then compiles these identified ranges into a final list of clustered sub-trajectories, resulting in a set of clusters that represent similar segments within the overall trajectory data.
@@
1 for i from 1 to $|C_{hist}|$+W-1 by step S do
2 $W_{s}\leftarrow i;$ $W_{e}\leftarrow i+W-1;$
3 for j from 1 to $|\Pi|$ do